For the pair of functions, find the quotient (f/g)(x) and give any x-values that are not in the domain of the quotient function.
F(x) = 14x^2-2 , g(x) 2x
Could you explain the steps if possible. Thank you

Answer by nerdybill(7384)

For the pair of functions, find the quotient (f/g)(x) and give any x-values that are not in the domain of the quotient function.
f(x) = 14x^2-2 , g(x) 2x
(f/g)(x) is simply f(x) divided by g(x):
(f/g)(x) = (14x^2-2)/(2x)
(f/g)(x) = 2(7x^2-1)/(2x)
(f/g)(x) = (7x^2-1)/x (this is the "quotient")
x-values that are NOT in the domain is only one number: 0
x cannot be zero, otherwise you would be dividing by zero (illegal)